Consider a normal distribution with mean 39 and standard deviation 7. What is the probability a value selected at random from this distribution is greater than 39? (Round your answer to two decimal places.)
Answer #1

Solution :

Given that ,

mean = = 39

standard deviation = = 7

P(x > 39) = 1 - P(x < 39)

= 1 - P((x - ) / < (39 - 39) / 7)

= 1 - P(z < 0)

= 1 - 0.50   

= 0.50

Probability = 0.50